OpenStudy (anonymous):

f(x)=11x+7 f-1(x)=__________ can anyone help me here with finding the inverse step by step?

OpenStudy (anonymous):

make it x = 11y + 7 and solve for x

OpenStudy (anonymous):

solve for y sorry

OpenStudy (anonymous):

so you should get (x-7)/11

OpenStudy (anonymous):

so (x-7)/11 is my answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you a bunch, hey can you help me solve a few others? but can you show me step by step how you are getting the answer?

OpenStudy (anonymous):

yea go ahead

OpenStudy (anonymous):

i'll redo the first one...instead of f(x) just make it x....then replace the x's in the equations with y...so you get x = 11y+7 step 1. move the 7 to the other side....x-7 = 11y. step 2. isolate your variable by dividing by 11. so you get y = x-7/11

OpenStudy (anonymous):

f(x)=3-4x find f-1(4)

OpenStudy (anonymous):

x = 3-4y......x-3 = -4y ........x-3/ -4 is the inverse...but now they want to know what f -1 (4) is so you plug 4 for x into the inverse equation so it 4-3/-4 which is -1/4 as your answer

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

this is trickier, If f(x)=x+7 and g(x)=x-7 (a) f(g(x))= (b) g(f(x)= (c) Thus g(x) is called an_________function of f(x)?

OpenStudy (anonymous):

A) rewrite this as f(x) times g(x) b/c that's really what it's asking. so it's (x+7)(x-7) and now you just foil it out. so it's x^2 -7x+7x-49 and the 7x's go away b/c -7x +7x is 0....so your answer for A is x^2-49

OpenStudy (anonymous):

for B. it's just the opposite. it's g(x) times f(x) therefore it's now (x-7)(x+7) which will come out to the same exact answer. and when you see a binomial like x^2 - 49 it's known as the difference of squares since 7^2 is 49. so when you see something like x^2 -25....u can just see it's (x-5)(x+5). So your answer for C is difference of squares but i guess their looking for inverse as the missing word? I honestly don't know about C.

OpenStudy (anonymous):

do i include parentheses for a and b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u mean like do you include parentheseis around x^2-49? that doesn't matter b/c it's just one term